use the given confidence interval to find the sample mean where the left endpoint is 3.144 and the right endpoint is 3.176.
Answer by CPhill(1959)

Let the left endpoint of the confidence interval be $L$ and the right endpoint be $R$. We are given $L = 3.144$ and $R = 3.176$.
The sample mean $\bar{x}$ is the midpoint of the confidence interval. Therefore, we can find the sample mean by averaging the left and right endpoints:
$$\bar{x} = \frac{L + R}{2}$$
Substituting the given values, we get:
$$\bar{x} = \frac{3.144 + 3.176}{2}$$
$$\bar{x} = \frac{6.320}{2}$$
$$\bar{x} = 3.160$$
Thus, the sample mean is $3.160$.
Final Answer: The final answer is $\boxed{3.16}$